# NIMA POST-PROCESS BANNER TO BE REMOVED AFTER FINAL ACCEPTANCE TCAD modeling of radiation-induced defects in 4H-SiC diodes

### v2

@@ 2 Materials and methods
Due to the p++-implant and metalization covering the active diode area, the proposed TCAD model considers only bulk defects, neglecting contributions of interface defects and oxide charges. Results of existing literature characterizing radiation-induced defects in 4H-SiC are subject to differing manufacturers and material quality, thus widely dispersed and sometimes contradictory. The ensuing large parameter space served as a starting point for the optimization of individual values to best fit experimental results from [5]. The investigated publications focus on neutron and electron-irradiated 4H-SiC since results compare very well [8, 9, 10]. Defects are either considered intrinsic or radiation-induced via a linear introduction rate (), types, and cross-sections are assumed constant. Due to the low dark current levels of 4H-SiC, reverse I-V measurements are dominated by electric noise and surface currents, restricting TCAD comparison to forward bias operation. I-V simulations have been conducted as transient in SDevice, using an AC Analysis in Single-Device Mode, simultaneously obtaining C-V data ( AC-frequency [5]). Obtained field constellations were saved in steps and used to simulate detection performance, utilizing the HeavyIon model to simulate transient particle hits corresponding to the -source used in [5].
@@ 3 Results
Table 1 summarizes the optimized defects of the TCAD irradiation model and their respective parameters, as well as publications best fitting the final obtained values. While the Z1,2 and EH6,7 defects are both considered major lifetime killers in 4H-SiC, the EH4 defect cluster [11] proved to be crucial as well. As discussions about the type of the EH6,7 defect continue, the presented model strongly suggests it to be of donor-type [12, 13]. The Boron-centers (B & D [14, 15]) are of intrinsic nature and, aside from slight improvements in C-V conformity, do not significantly affect results.
Figure 1 depicts measured and simulated I-V results in forward bias. The model accurately predicts the 4H-SiC device losing its rectifying properties with increasing irradiation, allowing for forward biasing up to high voltages [5, 10]. Absolute values are in good agreement with experiments, despite increasingly underestimating current levels with rising fluence, which may be attributed to surface currents omitted in the model. Although the exact origin of this is not yet fully understood, simulated trap occupation and field distributions suggest a charge separation provoked by holes and electrons diffusing from the implant and buffer layer, respectively. Trapped and concentrated close to their origins by corresponding defects, an electric field barrier is formed mid-device, suppressing current flow. The current level therein is mainly determined by the majority carrier trap (Z1,2 & EH4) concentration. At sufficient bias, defects on either side become fully occupied, increasing the respective carrier lifetimes and restoring conducting properties, as can be observed for the lowest fluence in Figure 1. Simulations suggest the deep-level centers EH6,7 and EH4 being the driving force behind this process, while the latter one reaches full occupation first and determines the rectification voltage. Not displayed are reverse I-V simulations, which show a slight increase in leakage current with irradiation fluence that remains below an order of magnitude increase. Although in agreement with other studies [7, 10], no further conclusion can be made, as simulated currents are multiple orders of magnitude below the experimental measurement limitations of around .

### v3

@@ 2 Materials and methods
Due to the p++-implant and metalization covering the active diode area, the proposed TCAD model considers only bulk defects, neglecting contributions of interface defects and oxide charges. Results of existing literature characterizing radiation-induced defects in 4H-SiC are subject to differing manufacturers and material quality, thus widely dispersed and sometimes contradictory. The ensuing large parameter space served as a starting point for the optimization of individual values to best fit experimental results from [5]. The investigated publications focus on neutron and electron-irradiated 4H-SiC since results compare very well [8, 9, 10]. Active defects in the simulation are either intrinsic (D & B) or radiation-induced (Z1,2, EH6,7 & EH4) via a linear introduction rate (). Defect types and cross-sections are assumed constant over the whole range of neutron fluences. Due to the low dark current levels of 4H-SiC, reverse I-V measurements are dominated by electric noise and surface currents, restricting TCAD comparison to forward bias operation. I-V simulations have been conducted in SDevice, using an AC Analysis in Single-Device Mode to simultaneously obtain C-V data ( AC-frequency [5]) at equal simulation settings. Obtained field constellations were saved in steps and used to simulate detection performance, utilizing the HeavyIon model to simulate the depth-dependent energy loss of -particles corresponding to the source used in [5]. Applied settings and physics models are described in [11, 12, 13].
@@ 3 Results
Table 1 summarizes the optimized defects of the TCAD irradiation model and their respective parameters, as well as publications best fitting the final obtained values. Aside from the widely known major lifetime killers in 4H-SiC (Z1,2 & EH6,7), the EH4 defect cluster proved to be crucial as well. As discussions about the type of the EH6,7 defect continue, the presented model strongly suggests it to be of donor-type [14, 15]. Aside from slight improvements in C-V conformity, the intrinsic Boron centers (B & D) do not affect results.
Figure 1 depicts measured and simulated I-V results in forward bias. The model accurately predicts the 4H-SiC device losing its rectifying properties with increasing irradiation, allowing for forward biasing up to high voltages [5, 10]. Absolute values are in good agreement with experiments, despite increasingly underestimating current levels with rising fluence, which may be attributed to surface currents omitted in the model. Simulations suggest a high concentration of trapped holes (EH6,7) near the top, and electrons (EH4) near the bottom due to their respective abundance within the p++ implant and the n+ buffer layer [6]. The resulting potential forms an electric field barrier that suppresses current flow and increases with the irradiation fluence. The current level in this confinement is mainly determined by the majority carrier trap (Z1,2 & EH4) concentration. At sufficient bias, defects on either side become fully occupied, increasing the respective carrier lifetimes and restoring conducting properties, as can be observed for the lowest fluence in Figure 1. The resulting rectification voltage is determined by the EH4 concentration, as it reaches full occupations earlier. Not displayed are reverse I-V simulations, which show a slight increase in leakage current with irradiation fluence that remains below an order of magnitude increase. Although in agreement with other studies [7, 10], no further conclusion can be made, as simulated currents are multiple orders of magnitude below the experimental measurement limitations of around .